Multiple choice

Suppose that the concentration of ammonia produced in the reaction is 0.7 M and the concentrations of nitrogen and hydrogen used are 0.8 M and 0.02 M respectively. Calculate the equilibrium constant for the reaction. Start with the expression that follows this equation:

N2 (g) + 3H2 (g)---> 2NH3 (g)

a. 0

b. 1 x 10^-6

c. 100,000

d. 1,000,000
Answer #1

At the equilibrium value of [NH3] = 2x = 0.7 M

i.e. x = 0.35 M

Therefore, the equilibrium value of [N2] = 0.8 - x = 0.45 M

And the equilibrium value of [H2] = 0.02 - 3x = -1.03 M (which is impossible)

The equilibrium constant (Kc) = [NH3]2/[N2][H2]3

Note: There is something wrong with the given values, please check the values properly and follow the approach shown above.